OpenStudy (anonymous):

Mr. Jackson lengthened the fence along the back of his yard. Before he added 110 feet, the fence was 265 feet long. He divided the fence into 5-foot sections. How many 5-foot sections does he have now?

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

@j-

OpenStudy (anonymous):

@J-bird

OpenStudy (anonymous):

hang on, someone is already asking for my help brb

OpenStudy (anonymous):

@gorv do you know it

OpenStudy (anonymous):

i got 10 ???

OpenStudy (gorv):

total=265+110=375 5 feet section=375/5=75

OpenStudy (anonymous):

Mr. Jackson lengthened the fence along the back of his yard. Before he added 110 feet, the fence was 265 feet long. He divided the fence into 5-foot sections. How many 5-foot sections does he have now? 75 22 53 375

OpenStudy (anonymous):

wehats the answer

OpenStudy (gorv):

i already solve d it bro

OpenStudy (gorv):

see above

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

ok mthx i got more

OpenStudy (anonymous):

its not d

OpenStudy (anonymous):

You got your answer :)

OpenStudy (anonymous):

yea it was a

OpenStudy (gorv):

loll 375/5 =75 see properly its 75

OpenStudy (anonymous):

om thx

OpenStudy (anonymous):

During a carpet sale, a carpet company sold $5,880 worth of carpet. If they sold 840 square yards of carpet, what was the average price per square yard? $6 $14 $5,040 $7

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

@J-bird

OpenStudy (anonymous):

someone help

OpenStudy (anonymous):

Hey Caleb I have to go now bye

OpenStudy (anonymous):

noooooooooooo

OpenStudy (anonymous):

Sorry I have a road trip and I need to help my mum

OpenStudy (anonymous):

oh nvm